# https://github.com/autonomousvision/sledge

### v1

@@ 4.2 Lane Graph Generation
Baselines. We select four diverse baselines. (1) VAE: we train a convolutional VAE with a 2D decoder head to generate RSIs. (2) RVAE: we sample from the decoder $\phi$ of our proposed RVAE. (3) HDMapGen: an autoregressive hierarchical graph neural network for lane graph generation [32], reimplemented for nuPlan. (4) DiT (RSI): similar to the concurrent DriveSceneGen [42], this is an image diffusion model that generates the RSI.
Metrics. Lane graph generation does not have established evaluation protocols. Therefore, we use a comprehensive set of seven metrics. (1) Route Length measures the mean and std of the longest valid ego vehicle route in the 64m×$\times$64m FOV for 1k generated graphs. The remaining six metrics are Frechet distances in different feature spaces measured using 50k generated and ground truth graphs. Two of these involve rasterizing lane graphs into the RSI, and using the penultimate feature vector of a ResNet-50 for the Frechet distance. (2) ImNet uses a ResNet trained on ImageNet, and (3) RVEnc uses the encoder $\pi$ of our RVAE. The remaining metrics are taken from [10, 32], based on graph features used in urban planning. They operate on all nodes of the generated lane graphs with degree $\neq$ 2, which are referred to in [32] as key points. (4) Connectivity: this uses the degrees of all key points. (5) Density: the number of key points in the 64m×$\times$64m FOV. (6) Reach: the number of valid paths found from key points to others. (7) Convenience: Lengths of all valid paths from all key points. The Frechet metrics are scaled by suitable powers of 10 for readability.
Results. Our results are shown in Table 2. All DiT variants generate more plausible layouts than the VAE, RVAE and HDMapGen, with significantly better metrics. For DiT-L, we observe higher visual fidelity in intersection areas when using the RLM representation instead of the RSI. In particular, the RLM-based models excel at creating coincident endpoints between connected lanes, which is crucial for smooth simulation. Scaling to DiT-XL provides further gains, leading to results competitive to the upper bound of replay. Among the metrics used in our study, we find the CNN Frechet distances do not value connected drivable lanes. In particular, the ImNet Frechet distance increases when moving from the RSI to RLM or from DiT-L to DiT-XL, despite visual evidence of reduced fidelity (see supplementary for larger sets of randomly generated samples from each model). Therefore, we focus on Reach in our subsequent experiments.
@@ 4.3 SLEDGE Simulation of PDM-Closed
Results. We show our results in Table 3. Importantly, we note that setting up an evaluation with SLEDGE only requires a 3GB download of our DiT-L checkpoint for the Lane & Agent mode, and an additional 1GB download of the nuPlan maps for the Lane $\rightarrow$ Agent mode, in contrast to the 2TB of logs needed to set up nuPlan. When using easy routes and traffic, we observe similar PFRs for both replay-based and generative simulation despite this large compression factor. For replay, we observe over a 4$\times$ rise in PFR from 0.06 to 0.26 when extending the route from 100m to 500m. These are primarily due to PDM-Closed’s inability to make lane changes or overtake slow vehicles, which are important planning behaviors that are not strongly penalized on current benchmarks like Val14 [12]. In all settings, switching to hard routes increases the number of turns, which is more challenging than straight driving and in turn deteriorates the planning performance. In the most challenging settings with hard routes and traffic, the PFR increases to over 40%. We visualize some snapshots of states from 100m SLEDGE simulations in Fig. 1. Our supplementary video contains clips from such reactive simulations, where we highlight the weaknesses of PDM-Closed that were not sufficiently penalized in prior benchmarks.
@@ 5 Conclusion
Limitations. Evaluating simulators is hard. We provide metrics for the lane graph generation sub-task, and preliminary experiments on testing rule-based planning, but using the simulator for other downstream tasks, such as reinforcement learning, will be important to showcase its full potential. For efficiency, we use a relatively small FOV and simulation radius, and a simplistic lane representation consisting of only the centerline (assuming constant lane widths). These issues could be alleviated by further scaling of our model and compute. However, like other diffusion models, the compute requirements of our approach are already high. We see value in improving the efficiency of SLEDGE through relevant techniques for accelerating diffusion models [28, 31, 40, 27].

### v2

@@ 4.2 Lane Graph Generation
Baselines. We select four diverse baselines. (1) VAE: we train a convolutional VAE with a 2D decoder head to generate RSIs. (2) RVAE: we sample from the decoder $\phi$ of our proposed autoencoder. (3) HDMapGen: an autoregressive hierarchical graph neural network for lane graph generation [31], reimplemented for nuPlan. (4) DiT (RSI): similar to the concurrent DriveSceneGen [41], this is an image diffusion model that generates the RSI.
Metrics. Lane graph generation does not have established evaluation protocols. Therefore, we use a comprehensive set of four metric types. (1) Route Length measures the mean and std of the longest valid ego vehicle route in the 64m×$\times$64m FOV for 1k generated graphs. We assess sample quality with (2) Precision and coverage with (3) Recall based on the improved generative metrics from [21]. To obtain fixed-sized representations needed to estimate precision and recall with a nearest neighbour classifier, we rasterize the lane graphs and consider the penultimate feature vectors of a ResNet-50. We use two versions of these metrics with ResNets trained on ImageNet and the encoder $\pi$ of our autoencoder. The remaining metrics are (4) Frechet distances taken from [8, 31], based on graph features used in urban planning. They operate on nodes of the generated lane graphs with degree $\neq$ 2, which are referred to in [31] as key points. Connectivity: this uses the degrees of all key points. Density: the number of key points in the 64m×$\times$64m FOV. Reach: the number of valid paths found from key points to others. Convenience: Lengths of all valid paths from all key points. The Frechet metrics are scaled by suitable powers of 10 for readability. Precision, Recall, and Frechet distances are measured using 50k generated and ground truth graphs.
Results. Our results are shown in Table 2. All DiT variants generate more plausible layouts, with significantly better metrics than HDMapGen or the VAEs. For DiT-L, we observe higher coverage and visual fidelity when using the RLM representation instead of the RSI. In particular, the RLM-based models excel at creating coincident endpoints in intersections between connected lanes, which is crucial for smooth simulation. Scaling to DiT-XL provides further gains across all metrics, demonstrating the effectiveness of increasing the model capacity and compute budget. Since ImageNet features may be misaligned to BEV graphs, and the RVEnc features might favor RLM diffusion models, we focus on the urban planning Frechet metrics (in particular, Reach) in our subsequent experiments.
@@ 4.3 SLEDGE Simulation of PDM-Closed
Results. We show our results in Table 3. Importantly, we note that setting up an evaluation with SLEDGE only requires a 3 GB download of our DiT-L checkpoint for the Lane & Agent mode, and an additional 1 GB download of the nuPlan maps for the Lane $\rightarrow$ Agent mode, in contrast to the 2 TB of logs needed to set up nuPlan. When using easy routes and traffic, we observe similar PFRs for both replay-based and generative simulation despite this large compression factor. For replay, we observe over a 4$\times$ rise in PFR from 0.06 to 0.26 when extending the route from 100m to 500m. These are primarily due to PDM-Closed’s inability to make lane changes or overtake slow vehicles, which are important planning behaviors that are not strongly penalized on current benchmarks like Val14 [10]. In all settings, switching to hard routes increases the number of turns, which is more challenging than straight driving and in turn deteriorates the planning performance. In the most challenging settings with hard routes and traffic, the PFR increases to over 40%.
@@ 5 Conclusion
Limitations. Evaluating simulators is hard. We provide metrics for the lane graph generation sub-task, and preliminary experiments on testing rule-based planning, but using the simulator for other downstream tasks, such as reinforcement learning, will be important to showcase its full potential. For efficiency and robustness, we use a relatively small FOV and simulation radius, a simplistic lane representation consisting of only the centerline (assuming constant lane widths), and rule-based traffic behavior with IDM. These issues could be alleviated by further scaling of our model and extensions for learned motion behavior [50, 41]. However, like other diffusion models, the compute requirements of our approach are already high. We see value in improving the efficiency of SLEDGE through relevant techniques for accelerating diffusion models [27, 30, 39, 26].